Present Value Years Interest Rat Future Value Equation
$181 5 $317
$335 17 $1,080
$48,000 13 $185,382
$40,353 30 $531,618

Solve for the unknown interest rate in each of the following (round to 2 decimals)   
Answer #1
Present Value($) Years Interest Rate% Future Value($)
Solution 1 181 5 11.86% 317
Solution 2 335 17 7.13% 1080
Solution 3 48000 13 10.95% 185382
Solution 4 40353 30 8.97% 531618
Formula used Future Value= Present value(1+rate%)^no. of years
Solution 1 317=181(1+r)^5
317/181=(1+r)^5
(1.7513812)^(1/5)=1+r
1.118603-1=r
.118603=r
11.86%=r
Solution 2 1080=335(1+r)^17
1080/335=(1+r)^17
(3.22388)^(1/17)=1+r
1.071284-1=r
0.071284=r
7.13%=r
Solution 3 185382=48000(1+r)^13
185382/48000=(1+r)^13
(3.862125)^(1/13)=1+r
1.1095336-1=r
0.1095336=r
10.95%=r
Solution 4 531618=40353(1+r)^30
531618/40353=(1+r)^30
(1.317418)^(1/30)=1+r
1.089743-1=r
0.089743=r
8.97%=r

Hope this solve all your queries.